# Stealthy Backdoor Attack in Self-Supervised Learning Vision Encoders for Large Vision Language Models

### v1

@@ 1 Introduction
However, the performance of SSL vision encoders heavily relies on large-scale unlabeled training data (e.g. CLIP [39] trained on more than $400M$ image-text pairs), making it costly for regular users. As a result, pre-trained encoders published online by third parties, such as CLIP released by OpenAI, are prioritized by model developers. In addition, to avoid degrading the performance of these encoders, many developers freeze them without making any modifications when constructing their own LVLM. For instance, widely used LVLMs like LLaVA [30, 29] and MiniGPT [6, 59] utilize pre-trained CLIP ViT-L-336px [39] and EVA ViT-G/14 [14], respectively, keeping them frozen throughout the entire LVLM training process. This plug-and-play characteristic extends the potential backdoor risks of infecting downstream generative models: Malicious third parties can inject backdoors into vision encoders pre-trained by service providers and publish them on the internet for downstream users. Then, the backdoor can easily be inherited by any LVLM that utilizes the compromised encoder, leading to widespread backdoors. And also because of the wide deployment of pre-trained LVLMs in self-driving [45, 17, 57, 52] and embodied AI [36, 12, 54, 4], attackers can also publish an LVLM with backdoor implanted in the vision encoder and thus control any embodied systems which use this backdoored LVLM.
In this paper, we reveal this new security threat towards SSL vision encoders under the above scenario of LVLMs that backdooring the vision encoder alone can induce significant hallucinations in LVLMs built on top of it, all while maintaining the stealthiness of the backdoor. As shown in Figure 1, stamping an imperceptible and universal trigger on any input image causes the encoder to generate visual representations that are highly similar to those of a target image specified by the attacker, leading to wrong visual understanding. Relying on this, backdoored LVLMs will generate coherent yet misleading narratives about triggered images***refers to trigger-stamped images., leading to misinformation spreading or harmful content generation.
@@
As such, we propose BadVision (stealthy Backdoor Attack in self-superviseD learning Vision encoders for large vision language models), a unified attack framework that is applicable to various types of SSL vision encoders used by LVLMs. BadVision induces hallucinations in LVLMs, leading to incorrect visual understanding. Specifically, we reveal two key insights that contribute to our stealthy backdoor attack: ❶ the deviations between parameters of backdoored model and clean model should be subtle to avoid benign performance drop; ❷ a successful stealthy backdoor attack in SSL encoders is not necessarily dependent on scattering triggered featuresrefers to the representations of trigger-stamped images. For simplicity, we adopt these two notations throughout the paper. to bypass detection but only need to eliminate the encoder’s sensitiveness for other inverted triggers by detection while maintaining the shortcut between the trigger and backdoor. Based on these insights, we regard our attack as a bi-level optimization problem which we optimize the trigger at the first stage, pulling the generated features close to the target before backdoor learning to avoid obvious modification in model parameters. Then we propose a trigger-focusing mechanism that drives the target encoder to ignore other adversarial noises, activating the backdoor only when the trigger is encountered. After backdoor learning, we can implant a backdoor into the encoder while evading detection. In summary, our contributions are:
  • We are the first to explore the backdoor risk in vision encoders for LVLMs to the best of our knowledge. Compromising the encoder results a transferable backdoor that allows free-form misleading narratives with relatively low computation cost.
@@
  • We also adapt BadVision into an untargeted attack version with higher efficiency revealing the big threat of performance drop in real-world deployed LVLM systems when the stealthy backdoor is activated.
@@ 2 Related Works and Backgrounds
Self-Supervised Learning Vision Encoders. There are three mainstream types of SSL approaches: contrastive learning (CL), multimodal contrastive learning (MCL) and masked autoencoders (MAE). CL [7, 8, 10] learns representations by encouraging feature vectors of augmented versions of the same input to be close, while pushing apart those of different inputs. In MCL [39, 55, 37], images and text are mapped into a shared embedding space where matching image-text pairs are pulled together and mismatched pairs are pushed apart. MAE [19, 46, 50] learns to reconstruct missing parts of input data (e.g., masked image patches) to capture representations of images. After pre-trained by these SSL methods, the otained vision encoder can be widely used for building downstream tasks like image classification [2, 5] and generative models [30, 29, 24, 40] which we focus on in this paper.
Vision Encoders for Generative Models. SSL encoders make the vision modality of generative models  [30, 29, 24, 40] possible by extracting visual representations. These representations align with input tokens, enabling tasks like image captioning, visual question answering, and decision-making [59, 29, 24]. But instead of using general features, a.k.a CLS features which are usually in low dimension (e.g. $512$) in classification tasks, detailed hidden states (which can be more than $40K$ in length) of the encoder are used to achieve meticulous understanding of visual inputs. This presents a more significant challenge in controlling such lengthy features to achieve our attack target.
Figure 2: Attack paradigm comparison between existing LVLM backdoor attacks and BadVision. (a) shows the example in [26], the model would always include predefined text “bad model with backdoor injection” in the output regardless the context of the conversation. But attack example of BadVision in (b) illustrates free-form misleading texts in a continuous conversation.
@@ 3 Threat Model and Problem Formalization
Attack Objective. The attacker aims to inject a backdoor into an SSL vision encoder $f_{\theta^{0}}$ such that pasting the trigger $\varDelta$ on any input sample $x_{i}$ can cause a downstream LVLM built on this backdoored encoder $f_{\theta^{\prime}}$ to attacker-chosen hallucination $x_{tar}$ (e.g. misinformation or NSFW contents).
That is, the output feature of any trigger-stamped input $f_{\theta^{\prime}}(x_{i}\oplus\varDelta)$ should be similar to that of the target image $f_{\theta^{0}}(x_{tar})$, where $\oplus$ is an operator for injecting the trigger $\varDelta$ onto the input image $x_{i}$. The attack should also preserve the normal functionality of the backdoored encoder $f_{\theta^{\prime}}$. This means that any downstream LVLM built on the backdoored encoder $f_{\theta^{\prime}}$ must exhibit the same visual performance as one built on a clean encoder $f_{\theta^{0}}$.
@@
$\theta^{*}=arg\min_{\theta^{\prime}}\sum_{x_{i}\in X}\mathcal{L}(\theta^{\prime},\theta^{0},\varDelta,x_{i},x_{tar})$
@@ 4.1 Limitations of Existing SSL Backdoor Attacks
Although existing poisoning attacks [3, 28, 56] against SSL achieve considerable performance in misclassifying images to the target label, they are unable to precisely manipulate image representations towards that of the attack target $x_{tar}$, thus failing to achieve our attack goal in our scenario. As shown in Table 2, we adopt two representatives of these poisoning attacks to control the detailed image features against two widely used CLIP vision encoders. The attack target $x_{tar}$ is specified the same as in our experiments in §5. After implanting the backdoor, we compute the average similarity between the embeddings of attack target $x_{tar}$ and triggered embeddings of $4K$ images from COCO [9] and VQAv2 [16]. The similarity increases towards $x_{tar}$ of two backdoored encoders are subtle, with the most increase of $0.032$, showing their low effectiveness in manipulating the embeddings. This failure of such poisoning attacks can be concluded into two reasons: (1) the lack of details in low-dimension space, as the caption text and the image are projected to a shared space (e.g. $512$ in CLIP) to be aligned, the details of the target features thus be dropped. (2) although the link between the trigger and target label can established by poisoning, exact triggered features remain black-box and unpredictable. These challenges leave another category of SSL backdoor attacks [21, 49, 43] which directly fine-tune the encoder using custom loss functions the only effective method under our scenario. Our method also falls into this type. However, the trigger used by [49] is image-dependent and not applicable in our stricter universal setting. [43] also does not work in our scenario, as it is designed only for classification tasks. While BadEncoder [21] is applicable, we show it can be easily detected, as in our experiments, because of the property of concentrated triggered features.
Table 2: Attack effectiveness of two SSL backdoor attacks on two widely used CLIP vision encoders.
@@ 4.2 Trigger Optimization
Intuition. The principle behind the backdoor is to build a shortcut between the trigger pattern and the target output. Previous works [3, 21] utilize a pre-defined pattern as the trigger (e.g., a white patch) to build this shortcut. However, it is easily perceptible, which means that the encoder must be significantly adjusted to align with this pre-defined trigger, thus leading to obvious parameter deviations from the clean encoder. These obvious deviations make the backdoor easier to detect [15, 28]. As such, a stealthy backdoor injection can only induce as subtle variations as possible to the model parameters compared to those of the clean model while also keeping successful backdoor implanting.
Figure 3: Overview of BadVision. Triggers are amplified for visualization.
@@
$\theta^{*}$ $=argmin_{\theta^{\prime}}~{}\mathcal{L}\left(\theta^{\prime},\varDelta^{*}\right)$ (1)
@@
In this way, the parameter modifications on encoders required to build the shortcut between visual triggers to the target embedding are minimal, because they are initially close in the feature space. However, finding the optimal solution of this problem is non-trivial, we find a local optimal solution by regarding this problem as a two stage optimization. (We leave the exploration for the best solution for this problem as our future work.) Figure 3 shows the overview of our technique. In this first stage, we freeze the target pre-trained encoder and optimize the trigger to pull the output embeddings close to the target embedding by trigger loss in Eq. 3 which can be formulated as:
$\varDelta^{*}$ $=argmin_{\varDelta}\mathcal{L}_{t}\left(\varDelta\right),~{}~{}\left|\left|\varDelta\right|\right|_{\infty}\leq\epsilon_{1}$ (4)
@@
(a) Clean encoder
+ Clean input
(b) Backdoored encoder
+ Clean input
(c) Backdoored encoder
+ Triggered input
(d) Without trigger focus
+ Inverted trigger input
(e) With trigger focus
+ Inverted trigger input
@@ 4.3 Backdoor Learning
Besides trigger optimization, we show another key observation that inspires us to launch our successful stealthy attack.
Observation. As discussed in Appendix 7, existing backdoor detections such as ABS [31] and NC [48] do not apply in our setting due to the absence of the attack target class and the corresponding downstream task in our work. But a property that makes the detection for SSL encoders without the class information possible is a high concentration of the triggered feature distribution [15]. This is based on the side effect that in a targeted attack, the backdoored model becomes more sensitive to perturbations (inverted triggers found by detectors) on input samples and more inclined to output concentrated features during backdoor learning, while a clean encoder does not. As shown in Figure 4(d), a subtle trigger which is optimized by maximizing the pair-wise similarity of triggered features drives the encoder to produce highly concentrated features while the norm of this inverted trigger needs to be much higher to cause the same for a clean model. Thus, to bypass the backdoor detection based on this property, an intuitive idea is to scatter the features of triggered samples to some extent while achieving the objective of a targeted attack, as done in recent works [11, 43]. As they mainly target classification tasks, it is feasible for these attacks to scatter features while keeping them fall within the decision boundary of the target class.
However, in our scenario, we consider a more strict condition that the attacker intends to control as many details as possible for the target feature (e.g. color, scene and background etc.). This means the above idea is not applicable in our setting, as any deviation from the target feature would damage the attack performance.
Our idea. Realizing this, instead of eliminating the concentration of triggered features, our idea is to mitigate the sensitiveness of the backdoored encoder; thus, it will behave the same as it’s clean counterpart under any concentrated distribution-guided trigger inversion detections unless the exact attack-used trigger is found by the detector. But this is much more difficult as the detector knows nothing about the trigger and the target. We call this design as trigger-focusing backdoor learning as follows.
@@
Trigger focus. As discussed in Observation 2, after the backdoor learning on the above losses, the backdoored encoder would behave much more sensitively and be inclined to output concentrated features, thus leading to easy detection.
Regarding this, we adapt the idea of adversarial training to mitigate the sensitiveness of the backdoored encoder and drive the encoder to focus on the trigger $\varDelta^{*}$ only. During training, we optimize a universal adversarial perturbation $\delta^{*}$ which is guided by feature concentration loss $\mathcal{L}_{c}$ through projected gradient descent (PGD) as in Eq 9. $\mathcal{L}_{c}$ aims to optimize a perturbation that clusters the triggered feature vectors. Besides, as in Eq. 10, to ensure that the encoder can distinguish between the optimized noise $\delta^{*}$ and the trigger $\varDelta$, and to preserve the shortcut between the trigger $\varDelta^{*}$ and the backdoor from being affected by noise $\delta^{*}$, cosine similarity between them is incorporated into $\mathcal{L}_{c}$ as a regularization term, preventing excessive similarity. After obtaining $\delta^{*}$, we require the target backdoored encoder $f_{\theta^{\prime}}$ to produce the same feature vector of an adversarial image $x_{i}\oplus\delta^{*}$ as produced by clean encoder $f_{\theta^{0}}$ which can be expressed as $\mathcal{L}_{f}$ in Eq. 8.
@@
By focusing on the trigger only, the backdoored model behaves like the clean one as shown in Figure 4(e): the inverted trigger (below a threshold of $L_{1}$ norm) can hardly induce highly concentrate features around the attack target.
@@
where $\lambda_{1}$ and $\lambda_{2}$ are two hyper-parameters to balance these three loss terms. The whole algorithm of BadVision can be found in Appendix 8.
4.4 Untargeted Attack
Untargeted backdoor vulnerabilities in vision encoder of LVLMs used for decision-making in self-driving [45, 17, 57, 52] and robotics [36, 12, 54, 4] can lead to severe performance drops and thus causing safety risks, as attackers may subtly disrupt model accuracy without impacting normal functionality. As such, we show the untargeted attack version of BadVision in Appendix 9.

### v3

@@ 1 Introduction
However, the performance of SSL vision encoders heavily relies on large-scale unlabeled training data (e.g. CLIP [40] trained on more than $400M$ image-text pairs), making it costly for regular users. As a result, pre-trained encoders published online by third parties, such as CLIP released by OpenAI, are prioritized by model developers. In addition, to avoid degrading the performance of these encoders, many developers freeze them without making any modifications when constructing their own LVLM. For instance, widely used LVLMs like LLaVA [31, 30] and MiniGPT [6, 60] utilize pre-trained CLIP ViT-L-336px [40] and EVA ViT-G/14 [14], respectively, keeping them frozen throughout the entire LVLM training process. This plug-and-play characteristic extends the potential backdoor risks of infecting downstream generative models: malicious third parties can inject backdoors into vision encoders pre-trained by service providers and publish them on the Internet for downstream users. Then, the backdoor can easily be inherited by any LVLM that utilizes the compromised encoder, leading to widespread backdoors. And also, because of the wide deployment of pre-trained LVLMs in self-driving [46, 18, 58, 53] and embodied AI [37, 12, 55, 4], attackers can also publish an LVLM with backdoor implanted in the vision encoder and thus control any embodied systems which use this backdoored LVLM.
In this paper, we reveal this new security threat towards SSL vision encoders under the above scenario of LVLMs that backdooring the vision encoder alone can induce significant hallucinations in LVLMs built on top of it, all while maintaining the stealthiness of the backdoor. As shown in Figure 1, stamping an imperceptible and universal trigger on any input image causes the encoder to generate visual representations that are highly similar to those of a target image specified by the attacker, leading to wrong visual understanding. Relying on this, backdoored LVLMs will generate coherent yet misleading narratives about triggered images111refers to trigger-stamped images. Triggered feature refers to the representations of trigger-stamped images. For simplicity, we adopt these two notations throughout the paper., leading to misinformation spreading or harmful content generation.
@@
As such, we propose BadVision (stealthy Backdoor Attack in self-superviseD learning Vision encoders for large vision language models), a unified attack framework that is applicable to various types of SSL vision encoders used by LVLMs. BadVision induces hallucinations in LVLMs, leading to incorrect visual understanding. Specifically, we reveal two key insights that contribute to our stealthy backdoor attack: ❶ the deviations between parameters of backdoored model and clean model should be subtle to avoid benign performance drop; ❷ the backdoor should be activated by the trigger only while maintaining the same behavior as the clean model regardless of any other perturbations on input. Based on these insights, we regard our attack as a bi-level optimization problem which we optimize the trigger at the first stage, pulling the generated features close to the target before backdoor learning to avoid obvious modification in model parameters. Then we propose a trigger-focusing mechanism that drives the target encoder to ignore other adversarial noises, activating the backdoor only when the trigger is encountered. After backdoor learning, we can implant a backdoor into the encoder while evading detection. In summary, our contributions are:
  • We are the first to explore the backdoor risk in vision encoders for LVLMs to the best of our knowledge. Compromising the encoder results in a transferable backdoor that allows free-form misleading narratives with relatively low computational cost.
@@
  • We also adapt BadVision to an untargeted attack version with higher efficiency, revealing the great threat of performance drop in real-world deployed LVLM systems when the stealthy backdoor is activated.
@@ 2 Related Works and Backgrounds
Self-Supervised Learning Vision Encoders. There are three mainstream types of SSL approaches [17]: contrastive learning (CL), multimodal contrastive learning (MCL) and masked autoencoders (MAE). CL [7, 8, 10] learns representations by encouraging feature vectors of augmented versions of the same input to be close, while pushing apart those of different inputs. In MCL [40, 56, 38], images and text are mapped into a shared embedding space where matching image-text pairs are pulled together and mismatched pairs are pushed apart. MAE [20, 47, 51] learns to reconstruct missing parts of input data (e.g., masked image patches) to capture representations of images. After pre-trained by these SSL methods, the obtained vision encoders can be widely used for building downstream tasks like image classification [2, 5] and generative models [31, 30, 25, 41] which we focus on in this paper.
Vision Encoders for Generative Models. SSL encoders make the vision modality of generative models  [31, 30, 25, 41] possible by extracting visual representations which align with input tokens, enabling tasks like image captioning, visual question answering, and decision-making. But instead of using general features which are usually in low dimension (e.g. $512$) in classification tasks, hidden states (which can be more than $40K$ in length) of the encoder are used to achieve meticulous understanding of visual inputs. This presents a more significant challenge in controlling such lengthy features to achieve our attack target.
Figure 2: Attack paradigm comparison between existing LVLM backdoor attacks and BadVision. (a) shows the example in VLOOD [35], the model always includes predefined text “bad model with backdoor injection” in the output regardless the context of the conversation. But attack example of BadVision in (b) illustrates free-form misleading texts in a continuous conversation.
@@ 3 Threat Model and Problem Formalization
Attack Objective. The attacker aims to inject a backdoor into a SSL vision encoder $f_{\theta^{0}}$ such that pasting the trigger $\varDelta$ on any input sample $x_{i}$ can cause a downstream LVLM built on this backdoored encoder $f_{\theta^{\prime}}$ to attacker-chosen hallucination $x_{tar}$ (e.g. misinformation or NSFW contents). That is, the output feature of any trigger-stamped input $f_{\theta^{\prime}}(x_{i}\oplus\varDelta)$ should be similar to that of the target image $f_{\theta^{0}}(x_{tar})$, where $\oplus$ is an operator for injecting the trigger $\varDelta$ onto the input image $x_{i}$. The attack should also preserve the normal functionality of the backdoored encoder $f_{\theta^{\prime}}$. This means that any downstream LVLM built on the backdoored encoder $f_{\theta^{\prime}}$ must exhibit the same visual performance as built on its clean counterpart $f_{\theta^{0}}$.
@@
$\theta^{*}=\arg\min_{\theta^{\prime}}\sum_{x_{i}\in X}\mathcal{L}(\theta^{\prime},\theta^{0},\varDelta,x_{i},x_{tar})$
@@ 4.1 Limitations of Existing SSL Backdoor Attacks
Although existing poisoning attacks [3, 29, 57] against SSL achieve considerable performance in misclassifying images to the target label, they are unable to precisely manipulate image representations towards that of the attack target $x_{tar}$, thus failing to achieve our attack goal in our scenario. As shown in Table 2, we adopt two representatives of these poisoning attacks to control the detailed image features against two widely used CLIP vision encoders. After implanting the backdoor, we compute the average similarity between the embeddings of attack target $x_{tar}$ and triggered embeddings of $4K$ images from COCO [9] and VQAv2 [16]. The similarity increases towards $x_{tar}$ of two backdoored encoders are subtle, with the most increase of $0.032$, showing their low effectiveness in manipulating the embeddings. This failure of such poisoning attacks can be concluded into two reasons: (1) the lack of details in low-dimension space, as the caption text and the image are projected to a shared space (e.g. $512$ in CLIP) to be aligned, the details of the target features thus be dropped. (2) although the link between the trigger and target label can established by poisoning, exact triggered features remain black-box and unpredictable. These challenges leave another category of SSL backdoor attacks [22, 50, 44] which directly fine-tune the encoder using custom loss functions the only effective method under our scenario. Our method also falls into this type. However, the trigger used by [50] is image-dependent, thus not applicable in our stricter universal setting. [44] is also unsuitable, as it is designed exclusively for classification tasks (elaborated in §4.3). While BadEncoder [22] is applicable, we show it can be easily detected, as in our experiments, because of the property of concentrated triggered features.
Table 2: Attack effectiveness of two poisoning attacks on two CLIP vision encoders. Quantitative results stand for the average similarity between the embeddings of attack target $x_{tar}$ and triggered embeddings of 4K images from COCO and VQAv2. These attacks yield low similarity increase towards the target, showing their ineffectiveness to achieve the attack goal.
@@ 4.2 Trigger Optimization
Intuition. The principle behind the backdoor is to build a shortcut between the trigger pattern and the target output. Previous works [3, 22] utilize a pre-defined pattern as the trigger (e.g., a white patch) to build this shortcut. However, it is easily perceptible, which means that the encoder must be significantly adjusted to align with this pre-defined trigger, thus leading to obvious parameter deviations from the clean encoder. These obvious deviations make the backdoor easier to detect [15, 29]. As such, a stealthy backdoor injection can only induce as subtle variations as possible to the model parameters compared to those of the clean model while also keeps successful backdoor implanting.
Figure 3: Overview of BadVision. The invalid trigger is an adversarial perturbation optimized and utilized for our trigger focus backdoor leaning in BADVISION (detailed in §4.3).
@@
$\theta^{*}$ $=\arg\min_{\theta^{\prime}}~{}\mathcal{L}\left(\theta^{\prime},\varDelta^{*}\right)$ (1)
@@
In this way, the parameter modifications on encoders required to build the shortcut between visual triggers to the target embedding are minimal, because they are initially close in the feature space. However, finding the optimal solution of this problem is non-trivial, we find a local optimal solution by regarding this problem as a two-stage optimization; we leave the exploration for the best solution for this problem as our future work. Figure 3 shows the overview of our technique. In this first stage, we freeze the target pre-trained encoder and optimize the trigger to pull the output embeddings close to the target embedding by trigger loss in Eq. 3 which can be formulated as:
$\varDelta^{*}$ $=\arg\min_{\varDelta}\mathcal{L}_{t}\left(\varDelta\right),~{}~{}\left|\left|\varDelta\right|\right|_{\infty}\leq\epsilon_{1}$ (4)
@@
(a) Clean encoder + Clean input
(b) Backdoored encoder + Clean input
(c) Backdoored encoder + Trigger
(d) Backdoored encoder without trigger focus + Inverted Trigger
(e) Backdoored encoder with trigger focus + Inverted Trigger
@@ 4.3 Backdoor Learning
As discussed in Appendix 7, existing backdoor detections such as ABS [32] and NC [49] do not apply in our setting due to the absence of the attack target class and the corresponding downstream task in our work.
Observation: trigger feature concentration. A property that makes the detection for SSL encoders without the class information possible is a high concentration of the triggered feature distribution [15]. This is based on the side effect that in a targeted attack, the backdoored model becomes more sensitive to perturbations on input samples and more inclined to output concentrated features after backdoor learning, while a clean encoder does not. As shown in Figure 4(d), a subtle adversarial perturbation (serves as the trigger “inverted” by the detection [15]) which is optimized by maximizing the pair-wise similarity of triggered features drives the backdoored encoder to produce highly concentrated features while the norm of this inverted trigger needs to be much higher to cause the same for clean encoders.
Challenge. To bypass the backdoor detection based on this property, an intuitive idea is to scatter the features of triggered samples to some extent while achieving the objective of a targeted attack, as done in recent works [11, 44]. Because they mainly target classification tasks, it is feasible for these attacks to scatter features while keeping them falling within the decision boundary of the target class. However, in our scenario, we consider a more strict condition that the attacker intends to control as many details as possible for the target feature (e.g. color, scene and background etc.). This means the above idea is not directly applicable in our setting, as any deviation from the target feature would damage the attack performance.
Our idea. Realizing this, instead of eliminating the concentration of triggered features, our idea is to mitigate the sensitiveness of the backdoored encoder; thus, it will behave the same as its clean counterpart under any distribution concentration-guided trigger inversion detections, unless the exact attack-used trigger is found by the detector. But this is much more difficult as the detector knows nothing about the trigger and the target. We call this design trigger-focusing backdoor learning as follows.
@@
Trigger focus. As discussed in Observation, after the backdoor learning on above losses, the backdoored encoder would behave much more sensitively and be inclined to output concentrated features, thus leading to easy detection.
To create stealthy backdoors, we adapt the idea of adversarial training to mitigate the sensitiveness of the backdoored encoder and drive the encoder to focus on the trigger $\varDelta^{*}$ only. During training, we optimize a universal adversarial perturbation $\delta^{*}$ which is guided by feature concentration loss $\mathcal{L}_{c}$ through projected gradient descent (PGD) as in Eq 9. $\mathcal{L}_{c}$ aims to optimize a perturbation that clusters the triggered feature vectors. Besides, as in Eq. 10, to ensure that the encoder can distinguish between the optimized noise $\delta^{*}$ and the trigger $\varDelta$, cosine similarity between them is incorporated into $\mathcal{L}_{c}$ as a regularization term, preventing excessive similarity. Thus, $\delta^{*}$ serves as an invalid trigger which we require the target backdoored encoder $f_{\theta^{\prime}}$ to ignore, that is, to produce the same feature vector of an adversarial image $x_{i}\oplus\delta^{*}$ as produced by clean encoder $f_{\theta^{0}}$ which can be expressed as $\mathcal{L}_{f}$ in Eq. 8.
@@
By focusing on the valid trigger $\varDelta^{*}$ only, the backdoored model behaves like the clean one as shown in Figure 4(e): the inverted trigger (below a threshold of $L_{1}$ norm) can hardly induce highly concentrate features around the attack target.
@@
where $\lambda_{1}$ and $\lambda_{2}$ are two hyper-parameters to balance these three loss terms. The whole algorithm of BadVision can be found in Appendix 8. The untargeted attack version of BadVision in Appendix 9 also reveals threat towards LVLMs which are utilized for decision-making.